OpenStudy (ria23):

What is the radius for the circle given by the equation x^2+(y-1)^2=12?

OpenStudy (tkhunny):

Eyeball problem. Just look at it and write down the answer. \(\sqrt{12}\)

OpenStudy (jhannybean):

\[(x-h)^2 +(y-k)^2 = r^2\]\[\text{center} = (h~,~k)\]\[\text{radius}=\sqrt{r^2}=r\]

OpenStudy (ria23):

In this case, would my h be 1?

OpenStudy (zale101):

(h,k) we don't know the numbers that indicates the center it just gives the variables, the only number we have is the radius. If it was \[ (x-1)^2+(y-k)^2= \sqrt{12}\] then h would be 1

OpenStudy (tkhunny):

Why do you care about the center of the circle. It asks only for the radius.

OpenStudy (ria23):

First, I wanna point out that I'm not out for the answer and second, I'm sorry but I am so confused. >~< would my equation end up being \[(x-1)^{2}+(x+1)^{2}=\sqrt12\]?

OpenStudy (jhannybean):

Its not a matter of caring, it's simply defining what the formula gives, and how it relates to the question given.

OpenStudy (jhannybean):

@Ria23 your equation does not change, but the question you posted was simply asking for the radius of the circle. Idk why you changed the format of your problem.

OpenStudy (jhannybean):

I was simply relating the format of the equation of the circle and your equation :\[\color{blue}{(x-h)^2} +\color{red}{(y-k)^2} = \color{violet}r^2\]\[\color{blue}{(x+0)^2}+\color{red}{(y-1)^2}=\color{violet}{12}\]

OpenStudy (jhannybean):

So when you equate \(r^2 = 12\) you get \(r= \pm\sqrt{12}\) but since the radius is always positive, you will have \(r = \sqrt{12}\)

OpenStudy (ria23):

Oh... Oh, I get it now. Thank yhu for explaining that for me. I wasn't sure what everyone meant...

OpenStudy (jhannybean):

No problem. The center is just centered at your point (h, k) which in your case would be (0,1). That was just extra information.

OpenStudy (ria23):

Can yhu explain why that would be 0? I thought that if yhu have a variable without a number in front it automatically became a 1...

OpenStudy (jhannybean):

It's not the number in front that matters. It is 0 because your original equation states : \(x^2+(y-1)^2=12\) So in order to fit THIS equation into the format of the equation of a circle, I set it up like my colorful post above. Therefore I found that the x-component for the center is 0.

OpenStudy (jhannybean):

Does that make sense?

OpenStudy (ria23):

Yes it does. c: Thank yhu again for all yhur help! ^.^

OpenStudy (jhannybean):

Woo! awesome :)